Check my work Check My Work button is now enabledItem 5Item 5 1.42 points Huron Company produces a commercial cleaning compound known as Zoom. The direct materials and direct labor standards for one unit of Zoom are given below: Standard Quantity or Hours Standard Price or Rate Standard Cost Direct materials 7.30 pounds $ 2.25 per pound $ 16.43 Direct labor 0.65 hours $ 6.50 per hour $ 4.22 During the most recent month, the following activity was recorded: 16,600.00 pounds of material were purchased at a cost of $2.05 per pound. All of the material purchased was used to produce 2,000 units of Zoom. 1,200 hours of direct labor time were recorded at a total labor cost of $11,400. Required: 1. Compute the materials price and quantity variances for the month. 2. Compute the labor rate and efficiency variances for the month. (For all requirements, Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance). Input all amounts as positive values. Round your intermediate calculations to the nearest whole dollar.) Garrison 16e Rechecks 2017-10-19
Answer #1

Requirement 1

Direct Material Price Variance is the difference between actual price paid for purchased materials and their standard cost. It helps to monitor the costs incurred to produce the goods. Direct Material Quantity Variance measures efficiency in the use of material by comparing standard material usage for actual production with actual material used, the difference is valued at standard cost

Direct Material Price Variance = (Standard Price - Actual Price) x Actual quantity

In the given case, all the materials purchased has been used for production

Material Price Variance = (2.25 – 2.05) x 16,600 = $ 3320 F

Material Quantity Variance = (Standard Quantity for actual production - Actual Quantity) standard price

= [7.30 x 2000 - 16600] x 2.25

= [14600 - 16600] x 2.25

= $ 4500 U

Requirement 2

Direct Labor rate variance indicates the actual cost of any change from standard labor rate of remuneration. Direct labor efficiency variance depicts how efficient the direct labor was in making the actual output that was produced by the direct labor

Labor Rate variance

= (standard rate - Actual rate) x Actual labor hour used

= [6.50 - (11400/1200)] x 1200

= (6.5-9.5) x 1200

= $3600 U

Labor Efficiency Variance

= (Standard hours for actual production - Actual hours) standard rate

= (0.65 x 2000 - 1200) x 6.5

= (1300 - 1200) x 6.5

= $650 F

kindly upvote